In a trapezium ABCD, AB ll DC and DC = 2AB. EF ll AB where E and F lie on BC and AD respectively such that BE /EC = 4/3. Diagonal DB intersects EF at G . Prove that 7EF = 11AB.

Grade:10

1 Answers

shaurya pratap
6 years ago
refer rd sharma
ignoree this-
Dear Yash
length of string = L
mass = M
angular speed (w) = 2/ pi    revolutions/sec  = 2*2 pi / pi = 4 rad/sec
tension (T) = m *v2 / r
 = m* r* w2
 = M * L * 42
 = 16 ML Newton
Regards
sps (askIITians forum expert)